OpenStudy (anonymous):

integral (x^2)/((4-x^2)^(3/2))dx

OpenStudy (loser66):

@surjithayer why does it take too long to type?? hehehe

OpenStudy (anonymous):

\[I=-\int\limits \frac{- x^2+4-4 }{\left( 4-x^2 \right)^{\frac{ 3 }{ 2 }} }dx\] \[=-\int\limits \frac{ 4-x^2 }{ (4-x^2)^{\frac{ 3 }{ 2 }} }dx+4 \int\limits \frac{ 1 }{ (4-x^2)^{\frac{ 3 }{ 2 }} }dx\] \[=-\int\limits \frac{ dx }{ \sqrt{4-x^2}}+4I1+c=-\sin^{-1} \left( \frac{ x }{ 2 } \right)+4I1+c\] for I1 put x=2sin theta and complete.

OpenStudy (loser66):

Woahh... ...excellent solution to me.

OpenStudy (anonymous):

sorry for late . i typed twice but some how i failed to post.

OpenStudy (anonymous):

thank you!